OpenStudy (anonymous):

find the distance between (-12,17) and (-6,9)

OpenStudy (anonymous):

a) \[\sqrt{14}\] b) 10 c) 100 d) 14

OpenStudy (anonymous):

so it will be (-12,17) and (-6,9) \[\sqrt(-6 + 12)^2 + (9 - 17)^2\]

OpenStudy (anonymous):

sorry about the first equation

OpenStudy (anonymous):

it will be \[\sqrt(6)^2 +(-8)^2 = \sqrt(36) + (64) = \sqrt(100) = 10\]

OpenStudy (anonymous):

so the answer is b

OpenStudy (anonymous):

the quadratic equation will be\[\sqrt(x2 -x1)^2 + (y2 - y1)^2\]